# Luminis Stellarum et Machina: Applications of Machine Learning in Light Curve Analysis

### v1

@@ 3.1 Supervised Learning
Figure 5: Overview of supervised learning algorithms. (a) RF: An ensemble of decision trees with majority voting or averaging. (b) MLP: A feedforward neural network with input, hidden, and output layers. (c) RNN: Processes sequential data using recurrent connections (denoted as red lines) between hidden layers. (d) LSTM: A specialized RNN with memory cells to capture long-term dependencies, utilizing recurrent connections (denoted as red lines) for information flow. (e) SVM: Identifies an optimal hyperplane with maximum margin for class separation. (f) KNN: Classifies data based on the majority vote of the $k$ nearest neighbors. (g) CNN: Extracts hierarchical features using convolutional and pooling layers, followed by fully connected layers for classification.
Random Forest (RF) is an ensemble learning method composed of multiple decision trees, where each tree is trained on a randomly selected subset of the training data and features Breiman (2001). This randomness reduces correlation between trees, enhancing generalization and robustness. Additionally, RF has relatively few hyper-parameters, making it an efficient and widely used ML model. Figure 5(a) illustrates a simple algorithm of RF.
Naive Bayes is a classification method based on Bayes’ theorem, assuming conditional independence of features given the target class Friedman et al. (1997). It estimates class probabilities using conditional probability formulas and assigns samples to the most probable class. Bayesian regression, in contrast, applies Bayesian inference to regression problems by modeling parameter distributions instead of single estimates, capturing uncertainty in predictions. Both methods leverage Bayesian principles, making them effective for probabilistic modeling in ML.
Artificial Neural Network (ANN) is a nonlinear, adaptive computational model designed to process information through a network of interconnected processing units Haykin (1994). Inspired by biological neural networks, ANNs simulate learning and decision-making by adjusting connections based on input data, making them highly effective for complex pattern recognition and predictive modeling in distributed environments.
Multilayer Perceptron (MLP) is a common type of ANN composed of multiple layers of fully connected neurons arranged hierarchically. It consists of an input layer, one or more hidden layers, and an output layer. Each neuron, or node, processes information using weighted connections and typically employs activation functions like sigmoid or ReLU to introduce non-linearity Pal & Mitra (1992). Figure 5(b) illustrates a typical MLP structure.
Recurrent Neural Network (RNN) is a type of ANN with recurrent connections, designed to model sequential data for recognition and prediction Bengio et al. (1994). It utilizes high-dimensional hidden states with nonlinear dynamics Sutskever et al. (2011), where each hidden state depends on its previous state Mikolov et al. (2014). This structure enables RNN to store, recall, and process complex temporal patterns over long durations, allowing it to map input sequences to output sequences and predict future time steps Salehinejad et al. (2017). Figure 5(c) illustrates a typical RNN structure.
Long Short-Term Memory (LSTM) is a specialized type of RNN designed to handle sequential data while overcoming the vanishing and exploding gradient problems. Unlike traditional RNNs, LSTMs use memory cells instead of standard hidden units, with gated mechanisms controlling input, output, and information flow. These gates help retain important features from previous time steps, enabling LSTMs to effectively capture long-term dependencies and improve sequence modeling Hochreiter & Schmidhuber (1997); Le et al. (2015); Gers et al. (2000). Figure 5(d) illustrates a typical LSTM structure.
Support Vector Machine (SVM) is a widely used SL algorithm applied in various astronomical tasks (e.g., Kovács & Szapudi (2015); Krakowski et al. (2016); Hartley et al. (2017); Hui et al. (2018); Ksoll et al. (2018); Pashchenko et al. (2018)). It identifies an optimal hyperplane in an $N$-dimensional space to separate classes. In two dimensions, this hyperplane is a line dividing the plane so that each class falls on a different side. The optimal hyperplane maximizes the margin—the distance between the plane and the closest data points, known as support vectors. Once determined, the hyperplane acts as a decision boundary for classifying new data. Figure 5(e) illustrates an SVM hyperplane for a linearly separable two-dimensional dataset.
$k$-Nearest Neighbors (KNN) is a non-parametric, instance-based learning algorithm used for classification and regression. Unlike most supervised methods that build predictive models, KNN directly stores training data and makes predictions by measuring similarity between data points. It calculates distances between a query point and all training samples, selecting the $k$ closest neighbors to determine the output Mucherino et al. (2009); Keller et al. (1985). As a lazy learner, KNN requires no explicit training phase but can be computationally expensive for large datasets due to pairwise distance calculations. Figure 5(f) illustrates a typical KNN structure.
Convolutional Neural Network (CNN) is a specialized DL architecture designed for processing structured grid-like data, such as images and time series. Unlike MLPs, which use fully connected layers throughout, CNNs employ convolutional layers that apply learnable filters to small regions of the input, capturing spatial hierarchies and local patterns. These layers reduce the number of parameters while preserving essential features, making CNNs efficient for feature extraction. Pooling layers further refine representations by reducing spatial dimensions, enhancing robustness to variations in input. Fully connected layers at the network’s end aggregate extracted features for final classification or regression tasks Li et al. (2021); O’shea & Nash (2015); Yamashita et al. (2018). Figure 5(g) illustrates a typical CNN structure.
@@ 3.2 Unsupervised Learning
Figure 6: Basic process of unsupervised learning.
@@
Figure 7: Overview of unsupervised learning algorithms. (a) K-means: Partitions data into $k$ clusters by minimizing the variance within each cluster. (b) HC: Hierarchical clustering builds a hierarchy of clusters. (c) PCA: Reduces dimensionality by projecting data onto orthogonal axes that maximize variance, with the first principal component shown. (d) t-SNE: Embeds high-dimensional data into a low-dimensional space while preserving local structures. (e) AE: An autoencoder consisting of an encoder and decoder, used to learn a low-dimensional latent representation of the input data. (f) SOM: A self-organizing map that projects high-dimensional data onto a low-dimensional grid while preserving topological properties.
K-means is a widely used and well-known clustering algorithm valued for its simplicity and efficiency. It groups data points into distinct clusters based on their similarities by minimizing the variance within each cluster. The algorithm begins by randomly initializing $K$ cluster centroids, then assigns each data point to the nearest centroid based on a chosen distance metric, typically Euclidean distance. After assignment, the centroids are updated by computing the mean of all points within each cluster. This process iterates until convergence, which occurs when the centroid positions stabilize or a predefined iteration limit is reached MacQueen (1967); Hartigan & Wong (1979); Sinaga & Yang (2020). Figure 7(a) illustrates how K-means works, showing the iterative process of centroid updates and cluster assignments.
@@
Hierarchical clustering (HC) is another popular clustering algorithm that aims to build a hierarchy of clusters Ward Jr (1963). There are two main types of HC: Agglomerative Hierarchical Clustering (AHC) and Divisive Hierarchical Clustering (DHC). AHC, also known as the ”bottom-up” approach, starts with each data point as an individual cluster and iteratively merges the closest clusters based on a chosen distance metric until it forms one cluster consisting of all data points. In contrast, DHC, referred to as the ”top-down” approach, begins with all data points grouped into a single cluster and recursively splits them into smaller clusters until each point becomes its own cluster Murtagh & Contreras (2012); Johnson (1967). Figure 7(b) illustrates AHC, showing the process of merging clusters in a ”bottom-up” manner.
@@
Principal Component Analysis (PCA) is a widely used dimensionality reduction technique that transforms high-dimensional data into a lower-dimensional representation while retaining as much relevant information as possible Wold et al. (1987). It achieves this by identifying and projecting the data onto a set of orthogonal directions, known as principal components, which capture the maximum variance in the dataset. This transformation not only reduces the number of features but also mitigates the curse of dimensionality, enhances computational efficiency, and improves data interpretability by focusing on its most informative aspects Abdi & Williams (2010); Jolliffe & Cadima (2016). Figure 7(c) demonstrates the process of PCA.
@@
t-Distributed Stochastic Neighbor Embedding (t-SNE) is another dimensionality reduction technique primarily used for visualization of high-dimensional data in two or three dimensions Van der Maaten & Hinton (2008). The algorithm represents each high-dimensional object as a two or three dimensional point, ensuring that similar objects are placed close together, while dissimilar objects are positioned farther apart with high probability. Unlike linear methods such as PCA, t-SNE captures non-linear relationships, making it particularly effective for revealing clusters and intricate patterns in complex datasets Van Der Maaten (2014); Kobak & Berens (2019). Figure 7(d) illustrates the t-SNE process.
@@
Autoencoder (AE) is an ANN designed to learn efficient low-dimensional representations of input data, commonly used for tasks like compression, dimensionality reduction, and visualization Yang & Li (2015); Gianniotis et al. (2015, 2016). It consists of two main components: an encoder and a decoder. The encoder compresses the input data into a lower-dimensional representation, often referred to as the latent space, while the decoder reconstructs the original data from this compressed form. During training, the network optimizes its weights by minimizing the reconstruction error, typically measured as the squared difference between the input and the reconstructed output. Once trained, the bottleneck layer of the encoder provides a compact representation of the data, enabling its use in reduced-dimensional spaces for analysis or visualization. Figure 7(e) shows the structure of an AE.
@@
Self-organizing map (SOM), also known as a Kohonen map Kohonen (1982), is an unsupervised ANN that creates a low-dimensional, typically two-dimensional, representation of high-dimensional input data. During training, the map self-organizes to closely match the topology of the input dataset, preserving its structure. In astronomy, SOMs have been applied for tasks such as semi-supervised classification, regression, clustering, visualization of complex datasets, and outlier detection (e.g., Armstrong et al. (2015, 2016); Meusinger et al. (2017); Rahmani et al. (2018)). Unlike traditional neural networks, where weights are used to transform input values through activation functions, the weights in SOMs represent the coordinates of the output neurons in the input data space. These weight vectors act as prototypes or templates, capturing the essential features of the input dataset. This unique structure allows SOMs to effectively map and visualize high-dimensional data in a more interpretable, lower-dimensional form. Figure 7(f) illustrates the SOM process.

### v2

@@ 4.1 Supervised Learning
Figure 6: Overview of SL algorithms. (a) RF: An ensemble of DTs with majority voting or averaging. (b) MLP: A feedforward neural network with input, hidden, and output layers. (c) RNN: Processes sequential data using recurrent connections (denoted as red lines) between hidden layers. (d) LSTM: A specialized RNN with memory cells to capture long-term dependencies, utilizing recurrent connections (denoted as red lines) for information flow. (e) SVM: Identifies an optimal hyperplane with maximum margin for class separation. (f) KNN: Classifies data based on the majority vote of the $k$ nearest neighbors. (g) CNN: Extracts hierarchical features using convolutional and pooling layers, followed by fully connected layers for classification.
Decision Trees (DTs) are hierarchical, non-parametric models that recursively partition the feature space by applying conditions on features at internal nodes Quinlan (1986). The model is defined by this tree structure, with nodes, branches, and leaves representing features, decisions, and outcomes, respectively. The learning algorithm trains the tree by selecting splits that minimize impurity measures such as entropy for classification or mean squared error (MSE) for regression:
$\text{Entropy}(R)=-\sum_{k=1}^{K}p_{k}\log_{2}p_{k},$
$\text{MSE}(R)=\frac{1}{|R|}\sum_{i\in R}(y_{i}-\bar{y}_{R})^{2},$
where $p_{k}$ is the probability of class $k$ in region $R$, and $\bar{y}_{R}$ is the mean of the target values in $R$. The algorithm also includes strategies to mitigate overfitting and bias, such as pruning or limiting tree depth, and handles class imbalance using cost-sensitive splitting or balanced sampling. DTs naturally quantify uncertainty via the class probability distribution at leaf nodes. DTs are used for star–galaxy separation Vasconcellos et al. (2011), detecting active objects Zhao & Zhang (2008), redshift estimation Suchkov et al. (2005), and classifying stars, galaxies, and AGNs Golob et al. (2021).
Random Forests (RFs) are ensembles of $B$ DTs, which together define the model Breiman (2001). Each tree partitions the feature space into regions corresponding to its leaf nodes, where a region $R$ contains all samples reaching that leaf. The learning algorithm trains each tree on bootstrap samples with random feature subsets. Predictions are then aggregated from all trees via majority voting for classification or averaging for regression:
$\hat{y}=\text{majority vote}\left(\{f_{b}(x)\}_{b=1}^{B}\right)\quad\text{or}\quad\hat{y}=\frac{1}{B}\sum_{b=1}^{B}f_{b}(x),$
Randomization reduces correlation among trees and enhances robustness. Uncertainty can be estimated using out-of-bag (OOB) predictions as an internal validation measure. Random feature selection mitigates bias and overfitting, while class imbalance is handled through balanced class weights. Figure 6(a) illustrates a typical RF architecture. RFs are widely used for supervised tasks such as variable star or supernova classification Jenkins et al. (2012); Kim & Bailer-Jones (2016); Gomez et al. (2020), and can also be applied in unsupervised settings to estimate object similarity Reis et al. (2018a, b).
Artificial Neural Networks (ANNs) are nonlinear, adaptive computational models composed of interconnected units called neurons Haykin (1994). The model is defined by the network architecture, specifying layers, neurons, and connectivity. Each neuron computes an output:
$y=\phi(w^{T}x+b),$
where $\phi$ is an activation function, $w$ is a vector of connection weights, $x$ is the input vector, and $b$ is a bias term. The learning algorithm, such as backpropagation, adjusts the weights to minimize prediction error. Inspired by biological neural networks, ANNs simulate learning and decision-making by adapting connections based on input data, making them highly effective for complex pattern recognition and predictive modeling in high-dimensional domains. Class imbalance is addressed using weighted loss functions, while uncertainty can be approximated via dropout or Bayesian layers. ANNs are used for classification of galaxy spectra Folkes et al. (1996), modeling Spectral Energy Distribution (SED) of galaxies Silva et al. (2011), probing the opacity of Universe to cosmic gamma rays Singh et al. (2022), and prediction of the atmospheric fundamental parameters from stellar spectra Azzam et al. (2021).
Multilayer Perceptrons (MLPs) are feedforward ANNs composed of an input layer, one or more fully connected hidden layers, and an output layer. The model is defined by the architecture specifying the number of layers, neurons per layer, and activation functions. Each hidden layer computes:
$\mathbf{h}_{l}=\sigma(\mathbf{W}_{l}\mathbf{h}_{l-1}+\mathbf{b}_{l}),$
where $\sigma$ is an activation function, typically the rectified linear unit (ReLU) Pal & Mitra (1992), $\mathbf{W}_{l}$ and $\mathbf{b}_{l}$ are the weight matrix and bias vector for layer $l$, and $\mathbf{h}_{0}=\mathbf{x}$ is the input vector. The learning algorithm, such as backpropagation, adjusts the weights to minimize loss functions such as cross-entropy (classification) or MSE (regression). Uncertainty can be modeled using dropout or Bayesian layers. Regularization techniques such as L2 weight decay reduce bias and prevent overfitting. Class imbalance is typically addressed by applying class-weighted loss functions or oversampling minority classes. Figure 6(b) illustrates a typical MLP architecture. MLPs are widely used in astronomy for tasks such as star–galaxy classification and distinguishing galaxies, stars, and quasars Andreon et al. (2000); Abd-elaziem & Soliman (2023); Zeraatgari et al. (2024).
Recurrent Neural Networks (RNNs) are neural networks with recurrent connections, defined by an architecture that specifies the number of layers, hidden units, and the recurrence structure for sequential data Bengio et al. (1994). The hidden state $h_{t}$ at time step $t$ is computed as:
$h_{t}=\sigma(W_{h}h_{t-1}+W_{x}x_{t}+b),$
where $x_{t}$ is the input at time $t$, $W_{h}$ and $W_{x}$ are weight matrices, $b$ is a bias vector, and $\sigma$ is an activation function (often $\tanh$ or ReLU). The learning algorithm, such as backpropagation through time (BPTT), adjusts the weights to capture temporal dependencies for sequence modeling and prediction Sutskever et al. (2011); Mikolov et al. (2014). Uncertainty is often estimated using techniques like dropout. Bias and instability due to vanishing or exploding gradients are mitigated by gradient clipping and batch normalization. Class imbalance is addressed through sequence-level weighting or synthetic time-series generation. Figure 6(c) illustrates a typical RNN architecture. RNNs are widely used in astronomy for time-series tasks, including supernova and variable star classification Charnock & Moss (2017); Naul et al. (2018), gravitational wave denoising Shen et al. (2019), periodic variable detection Tsang & Schultz (2019), and real-time transient discovery Möller & de Boissière (2020); Muthukrishna et al. (2019).
Convolutional Neural Networks (CNNs) are DL models designed for grid-like data such as images and time series Li et al. (2021); O’shea & Nash (2015); Yamashita et al. (2018). A CNN model consists of learnable convolutional filters $W$ applied to local regions $X$ of the input, extracting spatially hierarchical features:
$(W*X)_{i,j}=\sum_{m,n}W_{m,n}X_{i+m,j+n},$
Pooling layers reduce spatial dimensions, improving robustness to translations and distortions, while fully connected layers aggregate extracted features for final predictions. The corresponding learning algorithm adjusts the filter weights and biases via backpropagation and gradient descent. Uncertainty can be estimated using dropout or test-time augmentation (TTA). In TTA, multiple augmented versions of the input (e.g., rotations, flips, or shifts) are passed through the trained CNN at inference time, and the predictions are aggregated, for example via averaging or voting. The variability of predictions across these augmentations provides an empirical measure of uncertainty, while the aggregation improves robustness by reducing sensitivity to input transformations and noise. Bias is further reduced through data augmentation during training, which improves model invariance. Class imbalance is addressed using loss functions such as focal loss, which downweight well-classified examples. Figure 6(g) illustrates a typical CNN architecture. CNNs excel in astronomy tasks such as galaxy morphology classification Dieleman et al. (2015); Aniyan & Thorat (2017), variable star classification Hon et al. (2017), supernova identification Brunel et al. (2019); Pasquet et al. (2019), photometric redshift estimation Hoyle (2016), cosmological parameter inference Ntampaka et al. (2020), and strong gravitational lensing detection Lanusse et al. (2018).
Long Short-Term Memory (LSTM) networks are a specialized type of RNN, defined by an architecture that includes memory cells and gating mechanisms (input, forget, and output gates) to overcome vanishing and exploding gradient problems Hochreiter & Schmidhuber (1997); Le et al. (2015); Gers et al. (2000). The cell state $c_{t}$ is updated as:
$c_{t}=f_{t}\odot c_{t-1}+i_{t}\odot\tanh(W_{c}[h_{t-1},x_{t}]+b_{c}),$
where $\odot$ denotes element-wise multiplication. The learning algorithm, typically BPTT, adjusts the weights of the gates and hidden units to capture long-term dependencies in sequential data. Figure 6(d) illustrates a typical LSTM architecture. LSTMs are used for anomaly detection in LCs Zhang & Zou (2018), classification of solar radio spectra Xu et al. (2019), detection of core-collapse supernovae Iess et al. (2023), and solar flare prediction Sun et al. (2022).
Support Vector Machines (SVMs) are SL models that define a hyperplane separating classes in an $N$-dimensional space. The model is characterized by the weight vector $w$, bias $b$, and, for non-separable data, slack variables $\xi_{i}$. The learning algorithm then identifies the optimal hyperplane by solving the optimization problem:
$\min_{w,b,\xi}\frac{1}{2}\|w\|^{2}+C\sum_{i=1}^{N}\xi_{i},$
$\text{subject to}\quad y_{i}(w^{T}x_{i}+b)\geq 1-\xi_{i},\quad\xi_{i}\geq 0,\quad i=1,\dots,N,$
where $C$ is a regularization parameter controlling the trade-off between margin size and classification error. Nonlinear decision boundaries are handled via kernel functions such as the radial basis function (RBF) kernel. Uncertainty can be estimated using Platt scaling Platt et al. (1999), and class imbalance is addressed by adjusting class weights or using synthetic oversampling methods. Figure 6(e) illustrates a typical SVM separating hyperplane. SVMs are widely applied in astronomy for star-galaxy separation Kovács & Szapudi (2015), identification of galaxies Krakowski et al. (2016), classification of strong gravitational lenses Hartley et al. (2017), identification of pre-main-sequence stars Ksoll et al. (2018), and variable stars Pashchenko et al. (2018).
$k$-Nearest Neighbors (KNN) are non-parametric, instance-based models that define a prediction as a function of the $k$ closest training samples according to a distance metric, typically Euclidean distance Mucherino et al. (2009); Keller et al. (1985). For a query point $x$, the model output is given by:
$\hat{y}(x)$ $=\text{majority vote}\left(\{y_{i}\mid x_{i}\in\mathcal{N}_{k}(x)\}\right),$
$\mathcal{N}_{k}(x)$ $=\left\{\text{the }k\text{ nearest neighbors of }x\right\},$
The learning algorithm then determines the neighbors at prediction time using the chosen distance metric. As lazy learners, KNNs require no explicit training but can be computationally expensive for large datasets. Uncertainty in KNN predictions can be quantified using distance-weighted voting, where closer neighbors have higher influence on the predicted class, or by computing confidence scores based on the relative distances of neighbors from the query point. Feature scaling (normalization or standardization) is important to reduce bias in distance calculations. Class imbalance is addressed by adapting $k$ or weighting votes to emphasize minority classes. Figure 6(f) illustrates a typical KNN structure. KNN is widely applied in astronomy for classifying celestial objects Li et al. (2008), photometric redshift estimation Zhang et al. (2013), star–galaxy separation Machado et al. (2016a), and solar flare prediction Li et al. (2007).
@@ 4.2 Unsupervised Learning
Figure 7: Basic process of UL. Cylinder represents database, rectangles indicate data processing steps and rounded rectangle shows the final output.
@@
Figure 8: Overview of UL algorithms. (a) K-means: Partitions data into $k$ clusters by minimizing the variance within each cluster. (b) HC: Hierarchical clustering builds a hierarchy of clusters. (c) PCA: Reduces dimensionality by projecting data onto orthogonal axes that maximize variance, with the first principal component shown. (d) t-SNE: Embeds high-dimensional data into a low-dimensional space while preserving local structures. (e) AE: An autoencoder consisting of an encoder and decoder, used to learn a low-dimensional latent representation of the input data. (f) SOM: A self-organizing map that projects high-dimensional data onto a low-dimensional grid while preserving topological properties.
K-means is a widely used clustering model valued for its simplicity and efficiency. The model partitions data into $K$ clusters, each represented by a centroid $\mu_{k}$, and aims to minimize the within-cluster variance:
$\min_{\{C_{k}\}}\sum_{k=1}^{K}\sum_{x_{i}\in C_{k}}\|x_{i}-\mu_{k}\|^{2},$
The corresponding K-means learning algorithm initializes $K$ cluster centroids randomly, then iteratively assigns points to the nearest centroid and updates centroids until convergence MacQueen (1967); Hartigan & Wong (1979); Sinaga & Yang (2020). Uncertainty is addressed through soft clustering variants, while bias is mitigated via careful centroid initialization. K-means is sensitive to class imbalance but can be adapted using density-based weighting. Figure 8(a) illustrates how K-means works, showing the iterative process of centroid updates and cluster assignments. K-means is used for classifying stellar spectra Almeida & Prieto (2013), distinguishing stellar clusters based on chemical abundances Garcia-Dias et al. (2019), determining membership probabilities in open clusters El Aziz et al. (2016), and classifying eclipsing binary LCs Modak et al. (2020).
@@
Hierarchical clustering (HC) is a popular clustering model that builds a hierarchy of clusters Ward Jr (1963). There are two main types of HC: Agglomerative Hierarchical Clustering (AHC) and Divisive Hierarchical Clustering (DHC). AHC, also known as the ”bottom-up” approach, models each data point as an individual cluster and defines a merging criterion based on a chosen distance metric. DHC, or the ”top-down” approach, models all data points initially as a single cluster and defines a splitting criterion to recursively divide clusters until each point becomes its own cluster Murtagh & Contreras (2012); Johnson (1967). The HC learning algorithm builds the cluster hierarchy using linkage criteria:
$d(C_{i},C_{j})=\begin{cases}\min_{x\in C_{i},\,y\in C_{j}}d(x,y)&\text{(single linkage)},\\ \max_{x\in C_{i},\,y\in C_{j}}d(x,y)&\text{(complete linkage)},\end{cases}$
Cophenetic correlation quantifies uncertainty, and linkage choice affects bias. Figure 8(b) illustrates AHC, showing the process of merging clusters in a ”bottom-up” manner. Applications of HC include identifying galaxy cluster substructures Yu et al. (2015), detecting superclusters of galaxies Santiago-Bautista et al. (2020), exploring the hierarchical structure of open clusters Yu et al. (2020), and clustering astronomical spectra Tian et al. (2025).
@@
Principal Component Analysis (PCA) is a dimensionality reduction model that transforms high-dimensional data into a lower-dimensional representation while retaining as much relevant information as possible Wold et al. (1987). The model identifies orthogonal directions, called principal components, which capture the maximum variance in the dataset. Projecting data onto these components reduces the number of features, mitigates the curse of dimensionality, enhances computational efficiency, and improves interpretability Abdi & Williams (2010); Jolliffe & Cadima (2016).
The PCA learning algorithm computes the eigenvectors of the covariance matrix and projects data onto the top-$k$ eigenvectors:
$X^{\prime}=XV_{k},\quad\Sigma=X^{T}X,$
where $V_{k}$ contains the top-$k$ eigenvectors. Reconstruction error estimates uncertainty, and standardization prevents feature scale bias. Figure 8(c) demonstrates the PCA process. PCA is applied in astronomy for searching galaxy-scale strong lenses Joseph et al. (2014), exploring abundances in different environments Ting et al. (2012), and analyzing LCs of variable stars Deb & Singh (2009).
@@
t-Distributed Stochastic Neighbor Embedding (t-SNE) is a dimensionality reduction model designed to represent high-dimensional data in two or three dimensions while preserving local similarity relationships Van der Maaten & Hinton (2008). The model maps similar objects close together and dissimilar objects farther apart, capturing non-linear relationships and revealing clusters and intricate patterns in complex datasets Van Der Maaten (2014); Kobak & Berens (2019). The t-SNE learning algorithm minimizes the Kullback-Leibler divergence between probability distributions in the high- and low-dimensional spaces:
$KL(P\|Q)=\sum_{i}P(i)\log\frac{P(i)}{Q(i)},$
where the perplexity parameter balances local and global structure, affecting uncertainty. Careful tuning is required to avoid initialization bias. Figure 8(d) illustrates the t-SNE process. t-SNE is applied in astronomy to identify SED outliers Quispe-Huaynasi et al. (2025), reduce spectral information Traven et al. (2017), quantify similarity between planetary systems Alibert (2019), and separate Type-2 AGN from HII galaxies Zhang et al. (2020).
@@
Autoencoder (AE) is an ANN model designed to learn efficient low-dimensional representations of input data, commonly used for tasks such as compression, dimensionality reduction, and visualization Yang & Li (2015); Gianniotis et al. (2015, 2016). The model consists of two main components: an encoder, which compresses the input data into a lower-dimensional latent representation, and a decoder, which reconstructs the original data from this compressed form. The AE learning algorithm optimizes network weights by minimizing the reconstruction error, typically measured as the squared difference between the input and reconstructed output:
$\mathcal{L}=\|x-g_{\theta}(f_{\phi}(x))\|^{2},$
where $f_{\phi}$ encodes and $g_{\theta}$ decodes. Dropout layers estimate uncertainty, while regularization prevents overfitting. Figure 8(e) shows the structure of an AE. AEs are applied in astronomy to estimate gravitational wave parameters Gabbard et al. (2022), identify strong gravitational lenses Cheng et al. (2020), estimate atmospheric parameters Yang & Li (2015), and detect anomalies in cosmic ray variations Mandrikova & Mandrikova (2022).
@@
Self-organizing map (SOM), also known as a Kohonen map Kohonen (1982), is an unsupervised ANN model that creates a low-dimensional, typically two-dimensional, representation of high-dimensional input data. The model represents the output neurons with weight vectors that act as prototypes or templates, capturing the essential features of the input dataset. This structure allows SOMs to effectively map and visualize high-dimensional data in a more interpretable, lower-dimensional form while preserving the input topology. The SOM learning algorithm organizes the map by minimizing the quantization error:
$Q=\sum_{i}\|x_{i}-w_{c(i)}\|^{2}.$
where $w_{c(i)}$ is the winning neuron. Quantization error measures uncertainty, while neighborhood preservation reduces topological bias. Figure 8(f) illustrates the SOM process. SOMs are applied in astronomy to estimate galaxy photometric redshift probability density functions (P